OpenStudy (unklerhaukus):

f(x) = 3sin(x)-2cos(x)

OpenStudy (unklerhaukus):

\[=√(2^2+3^2)cos(x- \phi); \phi=arctan(2/3) \] \[=√13cos(x-arctan(2/3)\]

OpenStudy (unklerhaukus):

is this the correct simplified form?

OpenStudy (jamesj):

no, there's a negative sign outside

OpenStudy (unklerhaukus):

where did i miss the sign?

OpenStudy (jamesj):

f(x) = -(2 cos x - 3 sin x) so f(x) = - sqrt(13) cos(x - phi) where phi = arctan(-3/2)

OpenStudy (unklerhaukus):

so the arctangent stays in the solution .?

OpenStudy (jamesj):

sure

OpenStudy (unklerhaukus):

cool thanks

OpenStudy (jamesj):

just be sure of course you're using arctan the function with range (-pi/2,pi/2), vs. arctan the relation.

OpenStudy (unklerhaukus):

i am not sure ,